What is the mass of
       (a) 1 mole of nitrogen atoms?
       (b) 4 moles of aluminium atoms
       (c) 10 moles of sodium sulphite (Na2SO3)?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(a) Mass of 1 mole of nitrogen atoms = 14g
(b) 4 moles of aluminium atoms
Mass of 1 mole of aluminium atoms = 27 g
Mass of 4 moles of aluminium atoms = 27 × 4 = 108 g
(c) 10 moles of sodium sulphite (Na2SO3)
Mass of 1 mole of Na2SO3= 2 × 23 + 32 + 3 ×16
=46 + 32 + 48 = 126 g
Mass of 10 moles of Na2SO3= 126 × 10
= 1260 g

Thanks